# DPZV: Resource Efficient ZO Optimization For Differentially Private VFL

### v1

@@ 4 Methodology
1:  Input: Data ${\mathcal{D}}$, batch size $B$, learning rate $\eta_{m}$, total iteration $T$, smoothing parameter $\lambda_{m}>0$, clipping threshold $C>0$, privacy parameter $\sigma_{dp}$
2:  Output: Parameter $w_{0},w_{m}$ for all parties $m\in[M]$
3:  Initialize $w_{0},w_{m}$ and set $t,t_{m}=0$ for all parties
4:  for $t=1,\ldots,T$ do
5:     Sample ready-to-update client $m\in\{1,\ldots,M\}$
6:     Client $m$ samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}$ with corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$.
7:     Client $m$ computes perturbed local embeddings $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ according to (6).
8:      (Forward) Client transmits embeddings to server.
9:     for $k=1,\ldots,K$ do
10:        Server computes $\Delta_{m}^{t}$ according to (7) and (8).
11:        Server updates global server with $\Delta_{m}^{t}$ via $i)$ ZO Optimization or $ii)$ SGD.
12:         (Backward) Client receives $\Delta_{m}^{t}$.
13:        Client performs local update using (9).
14:     end for
15:     Client update local time stamp $t_{m}=t$.
16:  end for
4.1 DPZV Training Procedure
Each client $m$ operates independently, maintaining its own local communication round $t_{m}$, while the server tracks a global round $t$. Whenever the server receives information from a client, it increments $t$. Upon receiving an update from the server, the client synchronizes by setting $t_{m}=t$, capturing the latest state. This asynchronous mechanism allows clients to progress without waiting for stragglers, improving throughput and minimizing idle time.
The server maintains a copy of the latest local embeddings $\tilde{h}^{t}_{m,i}$ for all clients $m\in[M]$ and data samples $\xi_{i}\in{\mathcal{D}}$. Due to the asynchronous nature of the algorithm, these copies may be stale, as they do not always reflect the most up-to-date model parameters of the clients. Let $\tilde{t}_{m,i}$ denote the client time when the server last updated $\tilde{h}^{t}_{m,i}$. The delay at server communication round $t$ can then be expressed as $\tau_{m}^{t}=t_{m}-\tilde{t}_{m,i},$ where the delayed model parameters are defined as:
$\tilde{\chi}^{t}=\{w_{1}^{t_{1}-\tau_{1}^{t}},\dots,w_{M}^{t_{M}-\tau_{M}^{t}}\},\quad\tilde{{\bm{w}}}^{t}=\{w_{0}^{t},\tilde{\chi}^{t}\}.$ (5)
Local Embedding Perturbation: For each global iteration $t$, a client $m$ is activated, and it samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}\in{\mathcal{D}}$ and the corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$. To approximate gradients via zeroth-order finite differences, client $m$ computes two perturbed local embeddings for the mini-batch,
$\{h_{m,i}^{t_{m}+}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}+\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}},$
$\{h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}-\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ (6)
@@ 4.1 DPZV Training Procedure
Server side ZO computation: After the server receives the local embeddings $h_{m,i}^{t_{m}}$ from client $m$, it updates its embeddings copy and do computation. For each embedding pair $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}$, it computes the difference in loss function ${\mathcal{L}}$ caused by perturbation, divided by the smooth parameter $\lambda_{m}$. Specifically, we define [^1] We slightly abuse the notation, and define
$\tilde{f}(w_{0},h_{m,i}^{t_{m}\pm};y_{i})={\mathcal{L}}(w_{0},\tilde{h}^{t}_{1,i},\dots,h_{m,i}^{t_{m}\pm},\dots,\tilde{h}^{t}_{M,i};y_{i})$. where we treat $\tilde{f}$ as a function of the server parameter $w_{0}$ and the perturbed local embeddings.
$\delta_{m,i}^{t,t_{m}}=\frac{\tilde{f}(w_{0},h_{m,i}^{t_{m}+};y_{i})-\tilde{f}(w_{0},h_{m,i}^{t_{m}-};y_{i})}{\lambda_{m}},$ (7)
To ensure controllable privacy, the server then clips each $\delta_{m,i}^{t}$ by a threshold $C$ to bound sensitivity: $\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})=\min\{\delta_{m,i}^{t,t_{m}},C\}.$ It then samples noise $z_{m}^{t_{m}}$ from a Gaussian distribution $\mathcal{N}(0,\sigma_{dp}^{2})$. This noise is added to the mean of the per-sample-clipped updates, yielding a differentially private gradient-like quantity:
$\Delta_{m}^{t}=\frac{1}{B}\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})+z_{m}^{t}.$ (8)
The server then performs a backward communication and sends $\Delta_{m}^{t}$ to client $m$. The server then updates its global model through two possible operations: $i)$ ZO Optimization, ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}g(w_{0};\xi_{i})$ (defined in (2)); or $ii)$ stochastic gradient descent (SGD), ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}\nabla_{w_{0}}F({\bm{w}};\xi_{i})$, depending on the constraints on computation resources. The adopted ZO update methodology is explained in Sec. 3.2.
On the client side, upon receiving $\Delta_{m}^{t}$, client $m$ updates its local parameter ${\bm{w}}_{m}$ with learning rate $\eta_{m}$:
@@
where ${\bm{u}}_{m}^{t_{m}}$ is the same vector used for local perturbation.
Communication Efficient Asymmetric Update: By leveraging the low communication cost of ZO information, sever-client communication can be performed for a more convergence efficient design. As a result, the algorithm performs $K$ iterations of (7)-(9), where the server reuses the forward information from client $m$, resulting in an asymmetric communication protocol. This design leverages the advantage that multiple rounds of ZO updates incur significantly less overhead compared to a single round of first-order information exchange.
This asymmetric communication scheme maximizes the utility of forward information, which effectively reduces the per-round communication cost. Our method is well-suited for resource-constrained environments—such as edge devices with limited VRAM or computation power, while still maintaining robust performance and scalability in VFL settings. We summarize the pipeline above in Algorithm 1.
@@ 5 Convergence Analysis
Under an asynchronous update system, the probability of one client participating in one communication round is independent of other clients and satisfies:
(client $\mathbb{P}(\text{client }m\text{ uploading})=q_{m}.$
@@
Under 5.1-5.3. If we denote $q_{*}=\min_{m}q_{m}$, $d_{*}=\max_{m}d_{m}$ where $d_{m}$ represent the dimension of model parameters on device $m$, and let all step sizes satisfy: $\eta_{0}=\eta_{m}=\eta\leq\min\{\frac{1}{\sqrt{Td_{*}}},\frac{B}{4L(B+8d_{0})+8\gamma_{1}(2d_{m}+B)}\}$, let the smoothing parameter $\lambda$ satisfy: $\lambda\leq\frac{1}{Ld\sqrt{T}}$, and let the clipping level $C$ satisfy:
$C\geq\max\left\{0,\frac{1}{2}L\lambda d-\ell\sqrt{8\log(2\sqrt{2\pi})}\right\},$ (10)
@@
$\textstyle\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{{\bm{w}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]\leq\mathcal{O}\bigg{(}\frac{\mathbb{E}\left[F^{0}-F^{T}\right]\sqrt{d_{*}}}{\sqrt{T}}\textstyle+\frac{d_{*}}{T}+\frac{(\sigma_{s}^{2}/B+\sigma_{dp}^{2})\sqrt{d_{*}}}{\sqrt{T}}+\frac{C^{2}\sqrt{d_{*}}}{B\sqrt{T}}\bigg{)},$ (11)
The first term, $\mathcal{O}(\frac{\mathbb{E}\left[F^{0}-F^{T}\right]\sqrt{d_{*}}}{\sqrt{T}})$, is influenced by the model’s initialization, $F^{0}$. This term also enjoys the same rate as ZO optimization in the centralized case [17]. The second term $\mathcal{O}(\frac{d_{*}}{T})$ is a standard term for ZOO methods based on the usage of ZO estimator updates.
The third term, $\mathcal{O}(\frac{(\sigma_{s}^{2}/B+\sigma_{dp}^{2})\sqrt{d_{*}}}{\sqrt{T}})$, captures the impact of various noise sources in the learning system. Here, $\sigma_{s}^{2}/B$ represents the noise introduced by stochastic gradients, although increasing the batch size $B$ reduces the noise level towards convergence, it also increases memory requirements for update computations. Similarly, $\sigma_{dp}^{2}$ corresponds to the variance of the injected DP noise. While reducing $\sigma_{dp}^{2}$ improves utility, it comes at the cost of weakening privacy guarantees. Consequently, this term encapsulates the fundamental trade-off between model performance, computational cost, and privacy budget.
The fourth term, $\mathcal{O}(\frac{C^{2}\sqrt{d_{*}}}{B\sqrt{T}})$, highlights the role of the clipping operation in ZO optimization. A larger bounded sensitivity leads to greater fluctuations in the ZO information, potentially slowing convergence. The clipping level, however, isn’t the smaller, the better. As shown in (10), the selection of $C$ is lower bounded by the choice of the smoothing parameter $\lambda$, the smoothness level $L$, and Lipschitz constant $\ell$. This emphasizes the importance of carefully controlling the sensitivity level to ensure stable and efficient learning.
@@ 6 Privacy Analysis
6.1 Threat Model and Assumptions
While data are kept local in the VFL framework, the communication of local embeddings and backward gradient during the training process, as well as the sharing of model weights post training, can pose threat to sensitive information such as label and features [21]. In this section, we talk about how our algorithm protect privacy under these threat models. We consider two scenarios:
Honest-but-curious: In the honest-but-curious threat model, participants adhere strictly to the protocols of VFL without deviating from agreed procedures. However, they may attempt to infer private information from intermediate results exchanged during training.
Post-training adversarial: This model focuses on attacks conducted after the training process. It assumes the attacker lacks access to intermediate results. The adversary only has access to the trained model and aims to infer sensitive information by analyzing the model’s behavior or responses.
6.2 Differential Privacy Guarantee
@@
Under 5.1-5.3, suppose the privacy parameter $\sigma_{dp}$ is $\sigma_{dp}=\frac{2C\sqrt{T}}{D\mu},$ where $D$ denotes the volume of the dataset, $T$ defines total iterations, and $\mu>0$ controls the privacy level. The training process of Algorithm 1 is seen to be $(\epsilon,\delta(\epsilon))$-differential private for $\forall\epsilon>0$, where
$\delta(\epsilon)=\Phi(-\frac{\epsilon}{\mu}+\frac{\mu}{2})-e^{\epsilon}\Phi(-\frac{\epsilon}{\mu}-\frac{\mu}{2})$ (12)
The theorem provides privacy guarantee under the "honest-but-curious" threat model, where one or a few malicious clients try to do inference attacks by collecting information of the system. By providing only the differentially private ZO information $\Delta_{m}^{t}$, the algorithm protects for both label inference attacks [22] and feature inference attacks [23], because the attacker cannot differentiate a single datum in the dataset ${\mathcal{D}}$ and label set ${\bm{Y}}$. The differential privacy parameter $\mu$ is related to $(\epsilon,\delta)$ through the relationship defined in (12). Given any two of these parameters, the third can be determined by solving (12), allowing full flexibility in controlling the privacy level.
Although the forward embeddings $h_{m,i}$ are not protected by differential privacy and may be vulnerable to feature attacks from the server, existing attacks primarily rely on access to model parameters $w_{m}$ for model inversion [24, 25] or require backward gradients for gradient inversion [26]. Both attack vectors are mitigated under our differential privacy guarantees. A detailed discussion of our framework’s robustness against various attacks is provided in Appendix E, and the detailed proofs can be found in Appendix C
@@ 7 Experiments
Table 1: Communication cost comparison. We compare the communication cost (measured in kBs) of one client per communication round for each VFL methods.
Method Forward Backward Total
VAFL 65.54 65.54 131.07
ZOO-VFL 131.07 0.51 131.58
VFL-CZOFO 65.54 68.10 133.64
DPZV(Ours) 65.54 0.51 66.05
We conduct extensive experiments on a variety of VFL tasks. We compare DPZV against SotA VFL methods under different levels of privacy, and empirically show that our DPZV algorithm achieves superior utility and faster convergence even under strict privacy constraints, while requiring less computation and communication resources.
7.1 Datasets and Experiment Setups
Dataset: We consider four datasets: image dataset MNIST [27], CIFAR-10 [28], semantic dataset Amazon Review Polarity [29], and multi-view dataset ModelNet40 [30]. For each dataset, we conduct a grid search on learning rates and other hyperparameters. We run 100 epochs on each method and select the best validation model.
(a) MNIST
(b) ModelNet40
(c) CIFAR-10
@@ 7.1 Datasets and Experiment Setups
Figure 2: Test Accuracy of VFL Methods on image classification Tasks without DP constraints. DPZV outperforms first-order VFL methods on two datasets and surpasses all other ZO-based methods across all three datasets, showing both a higher accuracy and a faster convergence rate.
(a) MNIST
(b) ModelNet40
(c) CIFAR-10
Figure 3: Comparison with other VFL methods under the same communication cost. The asymmetric communication design results in DPZV outperforming all baselines.
@@
For MNIST, we use a two-layer CNN model, for VFL’s data partitioning, we split the images by row evenly into 7 sub-images and assign them to 7 clients. For CIFAR-10 we use a four-layer CNN model, and partition each image into 2×2, 4 patches of the same size for 4 clients. For ModelNet40, we use a ResNet-18 model, and partition each object into 12 different camera views and allocate them to 12 clients. For Amazon Reviews, we use a pre-trained BERT [31] model, and split the tokenized data input into 3 paragraphs of the same number of tokens and distributed them across 3 clients. For all four datasets, we use a fully connected model of two linear layers with ReLU activations as the server model.
@@
Baselines: We compare our algorithm against several SotA VFL methods: 1) VAFL [2] 2) ZOO-VFL [7], 3) VFL-CZOFO [9]. All methods assume that the server holds the labels, and concatenates the embeddings of clients as the input of the server. VAFL updates its model through first-order optimization in an asynchronous manner. We use VAFL as a first-order baseline of VFL method. Contrary to VAFL, ZOO-VFL and ZOFO both adopt ZO optimization in their training procedure. ZOO-VFL was the first attempt to combine ZOO with VFL. It simply substitutes the first order optimization by ZO optimization in common VFL methods, thus enjoying some benefits of ZOO. VFL-CZOFO uses a cascade hybrid optimization method. It computes the intermediate gradient via ZOO, while keeping the back propagation on both server and client.
While VAFL updates server and client once for every communication round, some first-order VFL methods adopts multiple local updates in one communication round [12, 3]. We note that while our DPZV adopts an asymmetric communication protocol, it is vertical to update protocols and can be easily modified to have multiple local update steps in one communication round. In this paper, we only focus on the same update behavior as VAFL and use VAFL as our first-order baseline. For each experiment, we set the total server-client update iterations of DPZV to $K=2$, all choices of step size of each method are shown in Appendix F.
Table 2: Accuracy under different privacy level. Our algorithm consistently outperforms all baselines, and the training performance barely decays when increasing the strictness of the DP guarantee. This shows the robustness and stability of our method against all DP levels.
Algorithm MNIST CIFAR-10 ModelNet40 Amazon Review
No DP $\epsilon=10$ $\epsilon=1$ No DP $\epsilon=10$ $\epsilon=1$ No DP $\epsilon=10$ $\epsilon=1$ No DP $\epsilon=10$ $\epsilon=1$
VAFL 91.61 82.14 43.11 43.74 37.19 36.51 80.34 71.59 66.41 85.76 69.70 60.30
ZOO-VFL 88.55 56.53 16.77 37.32 22.58 13.73 37.86 30.32 20.56 68.32 63.68 55.29
VFL-CZOFO 93.82 68.93 68.83 43.06 32.52 31.73 39.09 30.61 28.47 81.87 72.03 59.99
DPZV (ours) 96.53 95.36 94.47 46.78 42.63 42.45 79.64 83.77 83.51 84.42 81.82 80.30
7.2 DPZV Comparison to Baselines

### v2

@@ 1 Initialize $w_{0},w_{m}$ and set $t,t_{m}=0$ for all parties
Training Procedure. Each client $m$ maintaining its own local communication round $t_{m}$, while the server tracks a global round $t$. Whenever the server receives information from a client, it increments $t$. Upon receiving an update from the server, the client synchronizes by setting $t_{m}=t$, capturing the latest state. This asynchronous mechanism allows clients to progress without waiting for stragglers, improving throughput and minimizing idle time. The server maintains a copy of the latest local embeddings $\tilde{h}^{t}_{m,i}$ for all clients $m\in[M]$ and data samples $\xi_{i}\in{\mathcal{D}}$. Due to the asynchronous nature of the algorithm, these copies may be stale, as they do not always reflect the most up-to-date model parameters of the clients. Let $\tilde{t}_{m,i}$ denote the client time when the server last updated $\tilde{h}^{t}_{m,i}$. The delay at server communication round $t$ can then be expressed as $\tau_{m}^{t}=t_{m}-\tilde{t}_{m,i},$ where the delayed model parameters are defined as:
$\textstyle\tilde{\chi}^{t}=\{w_{1}^{t_{1}-\tau_{1}^{t}},\dots,w_{M}^{t_{M}-\tau_{M}^{t}}\},\quad\tilde{{\bm{w}}}^{t}=\{w_{0}^{t},\tilde{\chi}^{t}\}.$ (5)
Local Embedding Perturbation. For each global iteration $t$, a client $m$ is activated, and it samples a mini-batch ${\mathcal{B}}_{m}^{t_{m}}\in{\mathcal{D}}$ and the corresponding IDs ${\mathcal{I}}_{m}^{t_{m}}$. To approximate gradients via zeroth-order finite differences, client $m$ computes two perturbed local embeddings for the mini-batch,
$\textstyle\{h_{m,i}^{t_{m}+}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}+\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}},$
$\textstyle\{h_{m,i}^{t_{m}-}\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}=\{h(w_{m}^{t_{m}}-\lambda_{m}{\bm{u}}_{m}^{t_{m}};\xi_{m,i})\}_{i\in{\mathcal{I}}_{m}^{t_{m}}}$ (6)
@@
Server Side ZO Computation. After the server receives the local embeddings $h_{m,i}^{t_{m}}$ from client $m$, it updates its embeddings copy and do computation. For each embedding pair $\{h_{m,i}^{t_{m}+},h_{m,i}^{t_{m}-}\}$, it computes the difference in loss function ${\mathcal{L}}$ caused by perturbation, divided by the smooth parameter $\lambda_{m}$. Specifically, we define [^1] We slightly abuse the notation, and define $\tilde{f}(w_{0},h_{m,i}^{t_{m}\pm};y_{i})={\mathcal{L}}(w_{0},\tilde{h}^{t}_{1,i},\dots,h_{m,i}^{t_{m}\pm},\dots,\tilde{h}^{t}_{M,i};y_{i})$. where we treat $\tilde{f}$ as a function of the server parameter $w_{0}$ and the perturbed local embeddings.:
$\textstyle\delta_{m,i}^{t,t_{m}}=\frac{\tilde{f}(w_{0},h_{m,i}^{t_{m}+};y_{i})-\tilde{f}(w_{0},h_{m,i}^{t_{m}-};y_{i})}{\lambda_{m}},$ (7)
Scalar DP Noise Injection. To ensure controllable privacy, the server then clips each $\delta_{m,i}^{t}$ by a threshold $C$ to bound sensitivity: $\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})=\min\{\delta_{m,i}^{t,t_{m}},C\}.$ It then samples noise $z_{m}^{t_{m}}$ from a Gaussian distribution $\mathcal{N}(0,\sigma_{dp}^{2})$. This noise is added to the mean of the per-sample-clipped updates, yielding a differentially private gradient-like quantity:
$\textstyle\Delta_{m}^{t}=\frac{1}{B}\sum_{i\in{\mathcal{I}}_{m}^{t_{m}}}\mathrm{clip}_{C}(\delta_{m,i}^{t,t_{m}})+z_{m}^{t}.$ (8)
The server then performs a backward communication and sends $\Delta_{m}^{t}$ to client $m$. The server then updates its global model through two possible operations: $i)$ ZO Optimization, ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}g(w_{0};\xi_{i})$ (defined in (3)); or $ii)$ stochastic gradient descent (SGD), ${\bm{w}}_{0}\leftarrow{\bm{w}}_{0}-\eta_{0}\nabla_{w_{0}}F({\bm{w}};\xi_{i})$, depending on the constraints on computation resources. The adopted ZO update methodology is explained in Section 3. On the client side, upon receiving $\Delta_{m}^{t}$, client $m$ updates its local parameter ${\bm{w}}_{m}$ with learning rate $\eta_{m}$:
@@
where ${\bm{u}}_{m}^{t_{m}}$ is the same vector used for local perturbation. Our method is well-suited for resource-constrained environments, such as edge devices with limited VRAM or computation power, while still maintaining robust performance and scalability in VFL settings. We summarize the pipeline above in Algorithm 1.
@@ 5 Convergence Analysis
Under an asynchronous update system, the probability of one client participating in one communication round is independent of other clients and satisfies: (client $\mathbb{P}(\text{client }m\text{ uploading})=q_{m}.$
@@
Under 5.1-5.3. Define ${\mathcal{F}}=\mathbb{E}[F^{0}-F^{T}]$. If we denote $q_{*}=\min_{m}q_{m}$, $d_{*}=\max_{m}d_{m}$ where $d_{m}$ represent the dimension of model parameters on device $m$, and let all step sizes satisfy: $\eta_{0}=\eta_{m}=\eta\leq\min\{\frac{1}{\sqrt{Td_{*}}},\frac{B}{4L(B+8d_{0})+8\gamma_{1}(2d_{m}+B)}\}$, let the smoothing parameter $\lambda$ satisfy: $\lambda\leq\frac{1}{Ld\sqrt{T}}$, and let the clipping level $C$ satisfy: $C\geq\max\left\{0,\frac{1}{2}L\lambda d-\ell\sqrt{8\log(2\sqrt{2\pi})}\right\},$
@@
$\textstyle\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[\mbox{$\left\lVert\nabla_{{\bm{w}}}F({\bm{w}}^{t})\right\rVert$}^{2}\right]\leq\mathcal{O}\bigg{(}\frac{{\mathcal{F}}\sqrt{d_{*}}}{\sqrt{T}}\textstyle+\frac{d_{*}}{T}+\frac{(\sigma_{s}^{2}/B+\sigma_{dp}^{2})\sqrt{d_{*}}}{\sqrt{T}}+\frac{C^{2}\sqrt{d_{*}}}{(\exp(C^{2})-1)B\sqrt{T}}\bigg{)},$ (10)
Discussion. The first term is influenced by the model’s initialization, $F^{0}$. This term also enjoys the same rate as ZO optimization in the centralized case [14]. The second term $\mathcal{O}(\frac{d_{*}}{T})$ is a standard term for ZOO methods based on the usage of ZO estimator updates. The third term captures the impact of various noise sources in the learning system. Here, $\sigma_{s}^{2}/B$ represents the noise introduced by stochastic gradients, and $\sigma_{dp}^{2}$ corresponds to the variance of the injected DP noise. While reducing $\sigma_{dp}^{2}$ improves utility, it comes at the cost of weakening privacy guarantees. Consequently, this term encapsulates the fundamental trade-off between model performance, computational cost, and privacy budget. The fourth term, quantifies the impact of the gradient clipping operation on convergence. As $C$ increases, increases, this term diminishes, effectively recovering the non-private case in the limit. However, the noise variance $\sigma_{dp}^{2}$ also scales linearly with $C$ , which can impede overall convergence. This trade-off underscores the importance of carefully tuning the sensitivity level to balance privacy preservation and learning efficiency.
Based on the privacy analysis in Section 6, we further substitute the level of DP variance $\sigma_{dp}=\frac{2C\sqrt{T}}{D\mu}$ into this convergence rate. By selecting $T={\mathcal{O}}\left(\frac{D^{2}\mu^{2}\sqrt{d}}{C^{2}}\left({\mathcal{F}}+\frac{\sigma^{2}}{B}+\frac{C^{2}}{e^{C^{2}}-1}\right)\right),$ the overall convergence upper bound becomes ${\mathcal{O}}\left(\frac{\sqrt{d}}{D\mu}\right),$ which matches the known rate of DP-SGD for first-order methods [5]. This result underscores a key insight of our analysis: although zeroth-order optimization typically converges more slowly than its first-order counterpart in non-private settings, it can achieve the same convergence rate when subject to differential privacy constraints.
@@ 6 Privacy Analysis
While data are kept local in the VFL framework, the communication of local embeddings and backward gradient during the training process, as well as the sharing of model weights post training, can pose threat to sensitive information such as label and features [29]. In this section, we talk about how our algorithm protect privacy under these threat models. We consider two scenarios: 1) Honest-but-curious: In the honest-but-curious threat model, participants adhere strictly to the protocols of VFL without deviating from agreed procedures. However, they may attempt to infer private information from intermediate results exchanged during training. 2) Post-training adversarial: This model focuses on attacks conducted after the training process. It assumes the attacker lacks access to intermediate results. The adversary only has access to the trained model and aims to infer sensitive information by analyzing the model’s behavior or responses.
@@
Under 5.1-5.3, suppose the privacy parameter $\sigma_{dp}$ is $\sigma_{dp}=\frac{2C\sqrt{T}}{D\mu},$ where $D$ denotes the volume of the dataset, $T$ defines total iterations, and $\mu>0$ controls the privacy level. The training process of Algorithm  1 is seen to be $(\epsilon,\delta(\epsilon))$-differential private for $\forall\epsilon>0$, where
$\textstyle\delta(\epsilon)=\Phi(-\frac{\epsilon}{\mu}+\frac{\mu}{2})-e^{\epsilon}\Phi(-\frac{\epsilon}{\mu}-\frac{\mu}{2})$ (11)
Discussion. The theorem provides privacy guarantee under the "honest-but-curious" threat model, where one or a few malicious clients try to do inference attacks by collecting information of the system. By providing only the differentially private ZO information $\Delta_{m}^{t}$, the algorithm protects for both label inference attacks [13] and feature inference attacks [25], because the attacker cannot differentiate a single datum in the dataset ${\mathcal{D}}$ and label set ${\bm{Y}}$. The differential privacy parameter $\mu$ is related to $(\epsilon,\delta)$ through the relationship defined in (11). Given any two of these parameters, the third can be determined by solving (11), allowing full flexibility in controlling the privacy level.
Although the forward embeddings $h_{m,i}$ are not protected by differential privacy and may be vulnerable to feature attacks from the server, existing attacks primarily rely on access to model parameters $w_{m}$ for model inversion [16, 18] or require backward gradients for gradient inversion [19]. Both attack vectors are mitigated under our differential privacy guarantees. A detailed discussion of our framework’s robustness against various attacks is provided in Appendix E, and the detailed proofs can be found in Appendix C.
@@ 7 Experiments
Figure 2: Test Accuracy of VFL Methods on image classification tasks under DP constraints. $\delta$ is set to $1\times 10^{-3}$. DPZV outperforms first-order VFL methods on two datasets and surpasses all other ZO-based methods across all three datasets, showing both a higher accuracy and a faster convergence rate.
Figure 3: Privacy-Accuracy tradeoff across different datasets and algorithms. We use a constant level of $\delta=1\times 10^{-3}$ and vary $\epsilon$ to simulate different privacy levels. Our algorithm consistently outperforms baselines under tight privacy budget, showing a slower decay in performance than baselines as $\epsilon$ decreases.
@@
Setup. We consider four datasets: image dataset MNIST [7], CIFAR-10 [20], semantic dataset Amazon Review Polarity [27], and multi-view dataset ModelNet40 [34]. For each dataset, we conduct a grid search on learning rates and other hyperparameters. We run 100 epochs on each method and select the best validation model. We run each algorithm under three random seeds and compute the sample variance for the generosity of our result. For MNIST, we use a two-layer CNN model, for VFL’s data partitioning, we split the images by row evenly into 7 sub-images and assign them to 7 clients. For CIFAR-10 we use a four-layer CNN model, and partition each image into 2×2, 4 patches of the same size for 4 clients. For ModelNet40, we use a ResNet-18 model, and partition each object into 12 different camera views and allocate them to 12 clients. For Amazon Reviews, we use a pre-trained BERT [8] model, and split the tokenized data input into 3 paragraphs of the same number of tokens and distributed them across 3 clients. For all four datasets, we use a fully connected model of two linear layers with ReLU activations as the server model.
@@
We compare our algorithm against several SotA VFL methods: 1) VAFL [4] 2) ZOO-VFL [39], 3) VFL-CZOFO [33]. All methods assume that the server holds the labels, and concatenates the embeddings of clients as the input of the server. VAFL updates its model through first-order optimization in an asynchronous manner, and achieves DP by adding random noise to the output of each local embedding. We use VAFL as a first-order baseline of VFL method. Contrary to VAFL, ZOO-VFL and ZOFO both adopt ZO optimization in their training procedure. ZOO-VFL substitutes the first order optimization by ZO optimization in common VFL methods. VFL-CZOFO uses a cascade hybrid optimization method that computes the intermediate gradient via ZOO, while keeping the back propagation on both server and client. To enforce DP and compare all methods fairly, we follow the approach in VAFL [4], applying the same DP mechanism to both VFL-CZOFO and ZOO-VFL, which involves clipping the embeddings and adding calibrated vector noise. While VAFL updates server and client once for every communication round, some first-order VFL methods adopt multiple local updates in one communication round [24, 3]. In this paper, we only focus on the same update behavior as VAFL. In addition, model delay has been manually adjusted for all methods based on the per-batch computation time on a single client to simulate client heterogeneity and ensure a fair comparison.